OpenStudy (anonymous):

Calculate the total cost of the two items, including 8.25% sales tax. Item: $15 Item: $50 Please Help Me!

OpenStudy (anonymous):

@Hero @nincompoop @douglaswinslowcooper @Preetha @Loser66 @radar @

OpenStudy (loser66):

15+50 =?

OpenStudy (anonymous):

65

OpenStudy (loser66):

65*\(\Large \frac{8.25}{100}\) =?

OpenStudy (anonymous):

? = 536 1/4 Is That Correct?

OpenStudy (loser66):

5.36

OpenStudy (anonymous):

forgot the decimal, oops.

OpenStudy (loser66):

now, the cost is $65 the tax is $5.36 total is??

OpenStudy (anonymous):

Total Is = 70.36

OpenStudy (loser66):

Yup

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Is that for both of them?

OpenStudy (anonymous):

Nevermine.

OpenStudy (anonymous):

(15 + 50)(1.0825)=answer